OpenStudy (anonymous):

A wheel makes 11 1/8 revolutions per minute. If it rotates for 60 minutes, how many revolutions does it makes?

OpenStudy (noseboy908):

Let's start by making this a ratio. We're given that for every minute, the wheel rotates 11.125 times. Now, what does it tell us next? For how long does it rotate in the period in question?

OpenStudy (anonymous):

60 minutes

OpenStudy (noseboy908):

Yep. Now, keeping in mind that the ratio we just made is "1:11.125", what is the minutes in the ratio (the one) multiplied by to get 60?

OpenStudy (anonymous):

11?

OpenStudy (noseboy908):

No, think again. What is the one in the ratio multiplied by to get 60? What value times one equals 60?

OpenStudy (anonymous):

125

OpenStudy (noseboy908):

Let's test it. 125 x 1 = ?

OpenStudy (anonymous):

1 x 60

OpenStudy (noseboy908):

I'm confused, which question are you answering?

OpenStudy (anonymous):

i need the answer

OpenStudy (noseboy908):

And I'm helping you to it. What do you multiply 1 by to get 60?

OpenStudy (anonymous):

It turns 11 1/8 times a minute. How many times does it turn in 60 minutes?

OpenStudy (anonymous):

i dont know thats why i am asking

OpenStudy (noseboy908):

Then let me help you, and stop asking for the answer. What times one equals 60?

OpenStudy (mathmale):

@Studygeek8 : the more you're involved in finding problem solutions, the more you're going to understand, remember and use later. Unfortunately, you've given away part of your game: "i need the answer." Giving out free answers with little or no involvement from you is not the aim of OpenStudy. You're given this situation: "A wheel makes 11 1/8 revolutions per minute. If it rotates for 60 minutes, how many revolutions does it make? Why not write that RATE as\[\frac{ 11\frac{ 1 }{ 8 }~revolutions }{ 1~minute }...\]that says the same thing as all those words, but in mathematical terms. Those units of measurement are essential; it is really insufficient if you respond with "11," "125," or "60" and so such units. For how long does the wheel turn? Look for this time period in the original question. Now, multiply:\[\frac{ 11\frac{ 1 }{ 8 }~revolutions }{ 1~minute }*\frac{ how~many~minutes? }{ 1 }\]

OpenStudy (mathmale):

Replace that "how many minutes?" with the appropriate number (being certain to include the units of measurement, "minutes." The multiply! Which unit of measurement cancels here? What is your final answer:

OpenStudy (mathmale):

I note that you DID provide the proper number of minutes (60), but that you didn't include the necessary unit of measurement there.

OpenStudy (triciaal):

@Studygeek8